# Three-dimensional fracton topological orders with boundary Toeplitz braiding

### v2

@@ Abstract
In this paper, we theoretically study a class of 3D non-liquid states that show exotic boundary phenomena in the thermodynamical limit. More concretely, we focus on a class of 3D fracton topological orders formed via stacking 2D twisted $\mathbb{Z}_{N}$ topologically ordered layers along $z$-direction. Nearby layers are coupled while maintaining translation symmetry along $z$ direction. The effective field theory is given by infinite-component Chern-Simons theory, with an integer-valued symmetric block-tridiagonal Toeplitz $K$-matrix whose size is thermodynamically large. With open boundary conditions (OBC) along $z$, certain choice of $K$-matrices exhibits exotic boundary “Toeplitz braiding”, where the mutual braiding phase angle between two anyons at opposite boundaries oscillates and remains non-zero in the thermodynamic limit. In contrast, in trivial case, the mutual braiding phase angle decays exponentially to zero in the thermodynamical limit. As a necessary condition, this phenomenon requires the existence of boundary zero modes in the $K$-matrix spectrum under OBC. We categorize nontrivial $K$-matrices into two distinct types. Each type-I possesses two boundary zero modes, whereas each type-II possesses only one boundary zero mode. Interestingly, the integer-valued Hamiltonian matrix of the familiar 1D “Su-Schrieffer-Heeger model” can be used as a non-trivial $K$-matrix. Importantly, since large-gauge-invariance ensures integer quantized $K$-matrix entries, global symmetries are not needed to protect these zero modes. We also present numerical simulation as well as finite size scaling, further confirming the above analytical results. Motivated by the present field-theoretical work, it will be interesting to construct 3D lattice models for demonstrating Toeplitz braiding, which is left to future investigation.
@@ I Introduction
It is known that the effective theories of (2+1)D Abelian topological orders are $K$-matrix Chern-Simons theories [15], which provide a framework to understand a large class of topologically-ordered states, including multilayer fractional quantum Hall effects, spin liquids, and others. They also offer a systematic platform to classify symmetry-enriched topological (SET) and symmetry-protected topological (SPT) phases in (2+1)D, see, e.g., Refs. [16, 17, 18, 19, 20, 21, 22]. In the following, we assume all these (2+1)D topological orders reside in the spacetime with coordinates parameterized by $x$, $y$, $t$. So, the wavefunctions of these states are defined on an $xy$ plane. Next, one can stack each (2+1)D theory one by one along $z$ direction in order to form a multilayer quantum state. If the layer number tends to infinity, one ends up with an authentic (3+1)D quantum state. For example, one can start with the Laughlin state described by the Chern-Simons theory $\mathcal{L}=\frac{m}{4\pi}a_{\mu}\partial_{\nu}a_{\lambda}\epsilon^{\mu\nu\lambda}$, where $a_{\mu}$ is a $U(1)$ gauge field and $m\in\mathbb{Z}$. By stacking the Laughlin states and coupling the nearby layers, we obtain the effective theory for (3+1)D fractional quantum Hall states. The transport properties of the lateral surfaces of these states have been widely studied both theoretically and experimentally [23, 24, 25, 26, 27, 28]. The resulting low-energy effective theory of a multilayer fractional quantum Hall state is a multi-component Chern-Simons (mCS) theory, described by the Lagrangian $\mathcal{L}=\sum_{I=1}^{N}\frac{m}{4\pi}a_{\mu}^{I}\partial_{\nu}a_{\lambda}^{I}\epsilon^{\mu\nu\lambda}$, where $N$ refers to the number of layers. In this context, the superscript $I$, originally defined as the layer index, represents the location of the gauge field along $z$ direction. If $N$ is pushed to infinity, the resulting theory is an aforementioned infinite-component Chern-Simons, i.e., iCS field theory.
@@
The paper is organized as follows. In Sec. II, we carry on an analytic analysis on a concrete iCS field theory with Toeplitz braiding, whose $K$-matrix is, in its mathematical form, equivalent to the Hamiltonian of Su-Schrieffer-Heeger model [36]. In Sec. III, we propose a general framework for sorting $K$-matrices with boundary zero modes. Within this framework, we identify two distinct types of $K$-matrices that possess boundary zero modes, providing them as illustrative examples. The results are shown in Table 1. To demonstrate the correspondence between boundary zero modes of $K$-matrices and nonlocal braiding statistics in $z$-direction in these models, we conduct numerical computations in Sec. IV. This work is concluded in Sec. V.
Table 1: A class of Toeplitz $K$-matrices (Eq. (2)) with boundary zero modes, where the block $A=\begin{pmatrix}n&m\\ m&0\end{pmatrix}$ and the block $B=\begin{pmatrix}l_{1}&l_{2}\\ l_{3}&l_{4}\end{pmatrix}$. Type-I $K$-matrix possesses 2 boundary zero mode, whereas Type-II $K$-matrix possesses 1 boundary zero modes. It should be noted that the “requirement” column is not the sufficient condition for Toeplitz braiding.
@@ II Construction of iCS field theory and Toeplitz Braiding
$\Theta_{\mathbf{p},\mathbf{q}}=2\pi p^{I}(K^{-1})_{IJ}q^{J},\mod 2\pi.$ (3)
@@ II.2 Boundary zero modes and Toeplitz braiding: A concrete example
In the case of $|m|>|l|$, the mutual braiding phase angle between $\mathbf{l}_{I}$ and $\mathbf{l}_{J}$ is $2\pi[(K_{1}^{2N})^{-1}]_{IJ}$, decaying exponentially with respect to $|I-J|$ for $I$, $J$ satisfying $I>J,\ I=2i,\ J=2j-1,\ \text{or}\ J>I,\ J=2j,\ I=2i-1,\ i,j\in\mathbb{Z}^{+}$, which means that if two excitations reside at distinct $z$-boundaries, they cannot be mutually detected by braiding processes. For instance, the mutual braiding phase angle between two excitations $\mathbf{l}_{1}=\begin{pmatrix}1&0&\cdots&0\end{pmatrix}^{T}$ and $\mathbf{l}_{2N}=\begin{pmatrix}0&\cdots&0&1\end{pmatrix}^{T}$ is $\frac{2\pi}{m}\left(\frac{-l}{m}\right)^{N}\mod 2\pi$. Since $|l/m|<1$, the phase angle is suppressed to zero in thermodynamical limit $2N\rightarrow\infty$ as illustrated in Fig. 3, where we take $m=2,l=1$ as an example.
If $|l|>|m|$, the matrix element $[(K_{1}^{2N})^{-1}]_{IJ}$ increases exponentially with $|I-J|$ for $I$, $J$ satisfying $I>J,\ I=2i,\ J=2j-1,\ \text{or}\ J>I,\ J=2j,\ I=2i-1,\ i,j\in\mathbb{Z}^{+}$. The mutual braiding phase angle $\Theta_{\mathbf{l}_{1},\mathbf{l}_{2N}}=\frac{2\pi}{m}\left(\frac{-l}{m}\right)^{N}\mod 2\pi$ between $\mathbf{l}_{1}$ and $\mathbf{l}_{2N}$ is not exponentially suppressed if $l\mod m^{2}\neq 0$, which indicates that topological excitations $\mathbf{l}_{1}$ and $\mathbf{l}_{2N}$ can be remotely detected by each other in the braiding process. As the system size increases, the braiding phase angle between two excitations located at distinct $z$-boundaries, $\mathbf{l}_{1}$ and $\mathbf{l}_{2N}$, exhibits oscillation, as shown in Fig. 3.
@@
It is noteworthy that in some cases, the nonzero upper-right and lower-left elements may not result in nontrivial Toeplitz braiding. For example, if $l\mod m^{2}=0$, the upper-right and lower-left entries in Eq. (5) are integers, implying that excitations residing at distinct $z$-boundaries cannot interact by braiding. Indeed, the presence of boundary zero modes is a necessary but not sufficient condition for the appearance of Toeplitz braiding. We have demonstrated that the appearance of boundary zero modes in $K$-matrices will result in non-zero elements in the upper-right and lower-left regions of $K^{-1}$. However, the elements in these regions are not guaranteed to be non-integers. The $K_{1}$-type iCS field theory with $m=2$, $l=4$ is a typical example. The $K$-matrix has boundary zero modes, but nontrivial braiding statistics is completely absent except for the topological excitations residing at the same layer, as shown in Fig. 4(a). The presence of non-integers in the upper-right and lower-left regions of $K^{-1}$ is necessary for the occurrence of Toeplitz braiding. Therefore, we propose the following condition for the existence of Toeplitz braiding: If boundary zero modes of $K$ along with non-integer upper-right and lower-left regions of $K^{-1}$ exist, then Toeplitz braiding occurs, which is also illustrated in Fig. 4(b).
@@ III.3 Condition for Type-I and Type-II iCS field theories
$\left\{\begin{array}[]{l}|m|<|l_{2}+l_{3}|\\ l_{4}=0\\ n(l_{2}+l_{3})=2l_{1}m\\ l_{2}\neq l_{3}\end{array}\right.,$ (67)
@@
$\left\{\begin{array}[]{l}|m|<|l_{2}+l_{3}|\\ l_{4}=0\\ n(l_{2}+l_{3})\neq 2l_{1}m\\ l_{2}\neq l_{3}\end{array}\right.\text{ or }\left\{\begin{array}[]{l}|m|<|l_{2}-l_{3}|\\ l_{4}\neq 0\\ \det B=l_{1}l_{4}-l_{2}l_{3}=0\\ (2l_{2}m-l_{4}n)(2l_{3}m-l_{4}n)=0\end{array}\right.,$ (76)
@@
IV Numerical computation on Toeplitz braiding
@@
In this section, we conduct a numerical computation that illustrates the connection between the appearance of nonlocal braiding statistics in $z$-direction and the existence of boundary zero modes of $K$-matrix. For comparative analysis, we focus on iCS field theories that arise from stacking $\mathbb{Z}_{2}$ twisted topological orders. The block $A$ in the $K$-matrix, given by (2), is $\begin{pmatrix}2&2\\ 2&0\end{pmatrix}$ which describes the double-semion topological order.
@@ IV.1 Numerical computation on Toeplitz braiding via Type-I and Type-II iCS field theories
Fig. 5(d1-d3) are the matrix plots of the mutual braiding phase angle $\Theta_{\mathbf{l}_{I},\mathbf{l}_{J}}$ encoded in $K_{5}^{2N}$, where we also set the matrix size $402N=20,30,$ 40 during the numerical computation. The mutual braiding phase angle $\Theta_{\mathbf{l}_{I},\mathbf{l}_{J}}$ decreases with $|I-J|$, i.e., the distance in $z$ between topological excitations. If two topological excitations are sufficiently far apart in $z$-direction, they cannot detect each other by braiding. Fig. 5(d4) shows how the mutual braiding phase angle $\Theta_{\mathbf{l}_{1},l_{2N}}$ decreases drastically as the system size $2N$ increases, which also demonstrates that the topological excitations residing at distinct $z$-boundaries cannot detect each other via braiding.
@@ IV.2 Numerical computation on the relation between Toeplitz braiding and boundary zero modes
The matrix $M$ constructed by the boundary zero modes of $K_{2}^{2N}$, $K_{3}^{2N}$ and $K_{4}^{2N}$ are denoted as $M_{2}^{2N}$, $M_{3}^{2N}$ and $M_{4}^{2N}$, respectively. Compared to the inverse of the $K_{2}^{2N}$, $K_{3}^{2N}$ and $K_{4}^{2N}$, ${M_{2}^{2N}}$, $M_{3}^{2N}$ and $M_{4}^{2N}$ provide a good asymptotic approximation to the upper-right and lower-left elements, as depicted in Fig. 6(a1-a3), Fig. 6(b1-b3) and Fig. 6(c1-c3), where we define matrices $E_{j}^{2N}$ with element $(E_{j}^{2N})_{IJ}=|(M_{j}^{2N})_{IJ}-[(K_{j}^{2N})^{-1}]_{IJ}|$ ($j=2,3,4$) to characterize the difference between the two matrices $K_{j}^{-1}$ an $M_{j}$. To illustrate, we choose $402N=20,30,$ 40 for each type of iCS field theories. As $2N$ increases, the upper-right and lower-left elements of $M_{j}^{2N}$ and $(K_{j}^{2N})^{-1}$ ($j=2,3,4$) converge towards each other, as shown in Fig. 5(a4), Fig. 5(b4) and Fig. 5(c4), where we focus on the asymptotic behavior of $(E_{2}^{2N})_{2,2N}$, $(E_{3}^{2N})_{1,2N}$ and $(E_{4}^{2N})_{1,2N}$ as representatives. This observation suggests that we can extract information about the braiding statistics of excitations residing at distinct $z$-boundaries by analyzing the zero modes and their “close-to-zero” eigenvalues.

### v4

@@ Abstract
In this paper, we theoretically study a class of 3D non-liquid states that show exotic boundary phenomena in the thermodynamical limit. More concretely, we focus on a class of 3D fracton topological orders formed via stacking 2D twisted $\mathbb{Z}_{N}$ topologically ordered layers along $z$-direction. Nearby layers are coupled while maintaining translation symmetry along $z$ direction. The effective field theory is given by the infinite-component Chern-Simons (iCS) field theory, with an integer-valued symmetric block-tridiagonal Toeplitz $K$-matrix whose size is thermodynamically large. With open boundary conditions (OBC) along $z$, certain choice of $K$-matrices exhibits exotic boundary “Toeplitz braiding”, where the mutual braiding phase angle between two anyons at opposite boundaries oscillates and remains non-zero in the thermodynamic limit. In contrast, in trivial case, the mutual braiding phase angle decays exponentially to zero in the thermodynamical limit. As a necessary condition, this phenomenon requires the existence of boundary zero modes in the $K$-matrix spectrum under OBC. We categorize nontrivial $K$-matrices into two distinct types. Each type-I possesses two boundary zero modes, whereas each type-II possesses only one boundary zero mode. Interestingly, the integer-valued Hamiltonian matrix of the familiar 1D “Su-Schrieffer-Heeger model” can be used as a non-trivial $K$-matrix. Importantly, since large-gauge-invariance ensures integer quantized $K$-matrix entries, global symmetries are not needed to protect these zero modes. We also present numerical simulation as well as finite size scaling, further confirming the above analytical results. Symmetry fractionalization in iCS field theory is also briefly discussed. Motivated by the present field-theoretical work, it will be interesting to construct 3D lattice models for demonstrating Toeplitz braiding, which is left to future investigation.
@@ I Introduction
It is known that the effective theories of (2+1)D Abelian topological orders are $K$-matrix Chern-Simons theories [15, 16], which provide a framework to understand a large class of topologically-ordered states, including multilayer fractional quantum Hall effects, spin liquids, and others. They also offer a systematic platform to classify symmetry-enriched topological (SET) and symmetry-protected topological (SPT) phases in (2+1)D, see, e.g., Refs. [17, 18, 19, 20, 21, 22, 23]. In the following, we assume all these (2+1)D topological orders reside in the spacetime with coordinates parameterized by $x$, $y$, $t$. So, the wavefunctions of these states are defined on an $xy$ plane. Next, one can stack each (2+1)D theory one by one along $z$ direction in order to form a multilayer quantum state. If the layer number tends to infinity, one ends up with an authentic (3+1)D quantum state. For example, one can start with the Laughlin state described by the Chern-Simons theory $\mathcal{L}=\frac{m}{4\pi}a_{\mu}\partial_{\nu}a_{\lambda}\epsilon^{\mu\nu\lambda}$, where $a_{\mu}$ is a $U(1)$ gauge field and $m\in\mathbb{Z}$. By stacking the Laughlin states and coupling the nearby layers, we obtain the effective theory for (3+1)D fractional quantum Hall states. The transport properties of the lateral surfaces of these states have been widely studied both theoretically and experimentally [24, 25, 26, 27, 28, 29]. The resulting low-energy effective theory of a multilayer fractional quantum Hall state is a multi-component Chern-Simons (mCS) theory, described by the Lagrangian $\mathcal{L}=\sum_{I=1}^{N}\frac{m}{4\pi}a_{\mu}^{I}\partial_{\nu}a_{\lambda}^{I}\epsilon^{\mu\nu\lambda}$, where $N$ refers to the number of layers. In this context, the superscript $I$, originally defined as the layer index, represents the location of the gauge field along $z$ direction. If $N$ is pushed to infinity, the resulting theory is an aforementioned infinite-component Chern-Simons, i.e., iCS field theory.
@@
The paper is organized as follows. In Sec. II, we carry on an analytic analysis on a concrete iCS field theory with Toeplitz braiding, whose $K$-matrix is, in its mathematical form, equivalent to the Hamiltonian of Su-Schrieffer-Heeger model [37]. In Sec. III, we propose a general framework for sorting $K$-matrices with boundary zero modes. Within this framework, we identify two distinct types of $K$-matrices that possess boundary zero modes, providing them as illustrative examples. The results are shown in Table 1. To demonstrate the correspondence between boundary zero modes of $K$-matrices and nonlocal braiding statistics in $z$-direction in these models, we conduct numerical computations in Sec. IV. Numerical computations on symmetry fractionalization of iCS field theories with Toeplitz braiding are also briefly discussed in Sec. IV. This work is concluded in Sec. V.
Table 1: A class of Toeplitz $K$-matrices (Eq. (2)) with boundary zero modes, where the block $A=\begin{pmatrix}n&m\\ m&0\end{pmatrix}$ and the block $B=\begin{pmatrix}l_{1}&l_{2}\\ l_{3}&l_{4}\end{pmatrix}$. Type-I $K$-matrix possesses 2 boundary zero mode, whereas Type-II $K$-matrix possesses 1 boundary zero modes. Considering the physical meaning of block A, we require $m\neq 0$. It should be noted that the “requirement” column is not the sufficient condition for Toeplitz braiding.
@@ II Construction of iCS field theory and Toeplitz Braiding
$\Theta_{\mathbf{p},\mathbf{q}}=2\pi p^{I}(K^{-1})_{IJ}q^{J}\mod 2\pi.$ (3)
@@ II.2 Boundary zero modes and Toeplitz braiding: A concrete example
In the case $|m|>|l|$, the mutual braiding phase angle between $\mathbf{l}_{I}$ and $\mathbf{l}_{J}$ is $2\pi[(K_{1}^{2N})^{-1}]_{IJ}$, decaying exponentially with respect to $|I-J|$ for $I$, $J$ satisfying $I>J,\ I=2i,\ J=2j-1,\ \text{or}\ J>I,\ J=2j,\ I=2i-1,\ i,j\in\mathbb{Z}^{+}$, which means that if two excitations reside at distinct $z$-boundaries, they cannot be mutually detected by braiding processes. For instance, the mutual braiding phase angle between two excitations $\mathbf{l}_{1}=\begin{pmatrix}1&0&\cdots&0\end{pmatrix}^{T}$ and $\mathbf{l}_{2N}=\begin{pmatrix}0&\cdots&0&1\end{pmatrix}^{T}$ is $\frac{2\pi}{m}\left(\frac{-l}{m}\right)^{N-1}\mod 2\pi$. Since $|l/m|<1$, the phase angle is suppressed to zero in thermodynamical limit $2N\rightarrow\infty$ as illustrated in Fig. 3, where we take $m=2,l=1$ as an example.
If $|l|>|m|$, the matrix element $[(K_{1}^{2N})^{-1}]_{IJ}$ increases exponentially with $|I-J|$ for $I$, $J$ satisfying $I>J,\ I=2i,\ J=2j-1,\ \text{or}\ J>I,\ J=2j,\ I=2i-1,\ i,j\in\mathbb{Z}^{+}$. The mutual braiding phase angle $\Theta_{\mathbf{l}_{1},\mathbf{l}_{2N}}=\frac{2\pi}{m}\left(\frac{-l}{m}\right)^{N-1}\mod 2\pi$ between $\mathbf{l}_{1}$ and $\mathbf{l}_{2N}$ is not exponentially suppressed for $N$ being sufficiently large if $l$ is not divisible by $m~{}\mathrm{rad}(m)$ [^5] The radical $\mathrm{rad}(m)$ of an integer $m$ is the product of the distinct prime numbers dividing $m$., which indicates that topological excitations $\mathbf{l}_{1}$ and $\mathbf{l}_{2N}$ can be remotely detected by each other in the braiding process. As the system size increases, the braiding phase angle between two excitations located at distinct $z$-boundaries, $\mathbf{l}_{1}$ and $\mathbf{l}_{2N}$, exhibits oscillation, as shown in Fig. 3.
@@
It is noteworthy that in some cases, the nonzero upper-right and lower-left elements may not result in nontrivial Toeplitz braiding. For example, if $l$ is divisible by $m~{}\mathrm{rad}(m)$, the upper-right and lower-left elements in Eq. (5) are integers, implying that excitations residing at distinct $z$-boundaries cannot interact by braiding. Indeed, the presence of boundary zero modes is a necessary but not sufficient condition for the appearance of Toeplitz braiding. We have demonstrated that the appearance of boundary zero modes in $K$-matrices will result in non-zero elements in the upper-right and lower-left regions of $K^{-1}$. However, the elements in these regions are not guaranteed to be non-integers. The $K_{1}$-type iCS field theory with $m=2$, $l=4$ is a typical example. The $K$-matrix has boundary zero modes, but nontrivial braiding statistics is completely absent except for the topological excitations residing at the same layer, as shown in Fig. 4(a). The presence of non-integers in the upper-right and lower-left regions of $K^{-1}$ is necessary for the occurrence of Toeplitz braiding. Therefore, we propose the following condition for the existence of Toeplitz braiding: If boundary zero modes of $K$ along with non-integer upper-right and lower-left regions of $K^{-1}$ exist, then Toeplitz braiding occurs, which is also illustrated in Fig. 4(b).
@@ III.3 Condition for Type-I and Type-II iCS field theories
$\left\{\begin{array}[]{l}|m|<|l_{2}+l_{3}|,\\ l_{4}=0,\\ n(l_{2}+l_{3})=2l_{1}m,\\ l_{2}\neq l_{3},\end{array}\right.$ (67)
@@
$\left\{\begin{array}[]{l}|m|<|l_{2}+l_{3}|,\\ l_{4}=0,\\ n(l_{2}+l_{3})\neq 2l_{1}m,\\ l_{2}\neq l_{3},\end{array}\right.\text{ or }\left\{\begin{array}[]{l}|m|<|l_{2}-l_{3}|,\\ l_{4}\neq 0,\\ \det B=l_{1}l_{4}-l_{2}l_{3}=0,\\ (2l_{2}m-l_{4}n)(2l_{3}m-l_{4}n)=0,\end{array}\right.$ (76)
@@
In terms of the foliation structure[5, 9, 11], the Type-I and Type-II iCS field theories presented above are non-foliated. Within the context of iCS field theories, having a foliation structure means that one can apply $GL(N,\mathbb{Z})$ transformation $K\rightarrow K^{\prime}=WKW^{T}$ to extract a small block $K^{\prime}$, i.e. $K=K^{\prime}\oplus K^{\prime\prime}$, while leaving the Toeplitz structure of the matrix $K^{\prime\prime}$ unaltered. In order to preserve the locality in $z$-direction, $W\in GL(N,\mathbb{Z})$ is equal to the identity except for a small diagonal block, corresponding to a local relabeling of gauge fields. Based on the necessary condition established in Ref. [11], a necessary condition for the foliation structure is that the determinant polynomial
$D(u)$ $=\det(u^{-1}B+A+uB^{T})$
$=2l_{1}l_{4}-l_{2}^{2}-l_{3}^{2}-m^{2}+(l_{1}l_{4}-l_{2}l_{3})(\frac{1}{u^{2}}+u^{2})$
$+(l_{4}n-l_{2}m-l_{3}m)(\frac{1}{u}+u)$ (77)
must be a constant. For Type-I and Type-II iCS field theories, we can easily verify that $D(u)$ is not a constant. Therefore, Type-I and Type-II iCS field theories are non-foliated.
@@
IV Numerical computation on iCS Field Theories with Toeplitz braiding
@@
In Section IV.1 and IV.2, we conduct a numerical computation that illustrates the connection between the appearance of nonlocal braiding statistics in $z$-direction and the existence of boundary zero modes of $K$-matrix. For comparative analysis, we focus on iCS field theories that arise from stacking $\mathbb{Z}_{2}$ twisted topological orders. The block $A$ in the $K$-matrix, given by (2), is $\begin{pmatrix}2&2\\ 2&0\end{pmatrix}$, which describes the double-semion topological order. In Section IV.3, we present the symmetry fractionalization pattern in iCS field theories enriched by $U(1)$ charge conservation symmetry.
@@ IV.1 Numerical computation on Toeplitz braiding via Type-I and Type-II iCS field theories
Fig. 5(d1-d3) are the matrix plots of the mutual braiding phase angle $\Theta_{\mathbf{l}_{I},\mathbf{l}_{J}}$ encoded in $K_{5}^{2N}$, where we also set the matrix size $402N=20,30,$ 40 during the numerical computation. The mutual braiding phase angle $\Theta_{\mathbf{l}_{I},\mathbf{l}_{J}}$ decreases with $|I-J|$, i.e., the distance in $z$ between topological excitations. If two topological excitations are sufficiently far apart in $z$-direction, they cannot detect each other by braiding. Fig. 5(d4) shows how the mutual braiding phase angle $\Theta_{\mathbf{l}_{1},\mathbf{l}_{2N}}$ decreases drastically as the system size $2N$ increases, which also demonstrates that the topological excitations residing at distinct $z$-boundaries cannot detect each other via braiding.
@@ IV.2 Numerical computation on the relation between Toeplitz braiding and boundary zero modes
The matrix $M$ constructed by the boundary zero modes of $K_{2}^{2N}$, $K_{3}^{2N}$ and $K_{4}^{2N}$ are denoted as $M_{2}^{2N}$, $M_{3}^{2N}$ and $M_{4}^{2N}$, respectively. Compared to the inverse of the $K_{2}^{2N}$, $K_{3}^{2N}$ and $K_{4}^{2N}$, ${M_{2}^{2N}}$, $M_{3}^{2N}$ and $M_{4}^{2N}$ provide a good asymptotic approximation to the upper-right and lower-left elements, as depicted in Fig. 6(a1-a3), Fig. 6(b1-b3) and Fig. 6(c1-c3), where we define matrices $E_{j}^{2N}$ with element $(E_{j}^{2N})_{IJ}=|(M_{j}^{2N})_{IJ}-[(K_{j}^{2N})^{-1}]_{IJ}|$ ($j=2,3,4$) to characterize the difference between the two matrices $K_{j}^{-1}$ an $M_{j}$. To illustrate, we choose $402N=20,30,$ 40 for each type of iCS field theories. As $2N$ increases, the upper-right and lower-left elements of $M_{j}^{2N}$ and $(K_{j}^{2N})^{-1}$ ($j=2,3,4$) converge towards each other, as shown in Fig. 5(a4), Fig. 5(b4) and Fig. 5(c4), where we focus on the asymptotic behavior of $(E_{2}^{2N})_{2,2N}$, $(E_{3}^{2N})_{1,2N}$ and $(E_{4}^{2N})_{1,2N}$ as representatives.
This observation suggests that we can extract information about the braiding statistics of excitations residing at distinct $z$-boundaries by analyzing the zero modes and their “close-to-zero” eigenvalues.
Figure 7: (a1-a5) are the fractionalized charges $Q_{I}$ carried by topological excitations $\mathbf{l}_{I}$ in $K_{1,2,3,4,5}$-matrix iCS field theories with OBC applied in $z$-direction. In comparison, (b1-b5) are the fractionalized charges $Q_{I}$ of topological excitations with PBC applied in $z$-direction. For illustrative purpose, we select the system size $602N=$ 60.
IV.3 Charge fractionalization in iCS field theory
The discussion on Toeplitz braiding primarily addresses the impact of the boundary zero modes of $K$ on boundary physics. The influence of these boundary zero modes on $z$-bulk physics is manifested in the iCS field theory enriched by $U(1)$ charge-conservation symmetry. This subsection explores such symmetry enrichment and provides the results of numerical computation on charge fractionalization of topological excitations. Compared to the iCS field theories with periodic boundary condition (PBC) in the $z$-direction, we find that the pure mathematical entities-boundary zero modes of $K$, indeed influence the symmetry charges carried by topological excitations residing at $z$-bulk.
In the field theoretical formalism, symmetry charges are then carried by the currents $J^{I,\mu}=\frac{1}{2\pi}\epsilon^{\mu\nu\lambda}\partial_{\nu}a_{\lambda}^{I}$. In order to probe the conserved currents $J^{I,\mu}$, each $J^{I,\mu}$ is minimally coupled to an external electromagnetic field $A_{\mu}$. Consequently, in addition to the pure Chern-Simons term (Eq. (1)) and the excitation term, we incorporate an additional term $\sum_{I=1}^{2N}\frac{q^{I}}{2\pi}A_{\mu}\epsilon^{\mu\nu\lambda}\partial_{\nu}a_{\lambda}^{I}$ into the Lagrangian[16]. $q^{I}\in\mathbb{Z}$ composes a so-called charge vector $\mathbf{q}=(q^{1},\ldots,q^{2N})^{T}$. By integrating $a_{\mu}^{I}$ out, we obtain the effective $A_{\mu}$ charge $Q_{I}$ carried by the topological excitation $\mathbf{l}_{I}$ (Eq. (6)).
$Q_{I}=\mathbf{q}K^{-1}\mathbf{l}_{I}.$ (83)
Here, we select $q^{I}=1$ for $I=1,\ldots,2N$, which implies the trivial excitations carry integral $A_{\mu}$ charge. As a result, in the study of symmetry fractionalization, the key aspect is $Q_{I}\mod 1$, as the integral part of $Q_{I}$ does not matter. In the following discussion, we drop the integral part of $Q_{I}$. The effective $A_{\mu}$ charges $Q_{I}$ carried by $\mathbf{l}_{I}$ ($I=1,\ldots,2N$) in $K_{1}$(with $m=2$, $l=3$), $K_{2}$, $K_{3}$ and $K_{4}$-matrix iCS field theory are plotted in Fig. 7(a1-a4). For illustrative purposes, we set the system size $602N=$ 60. $Q_{I}$ exhibits random oscillations as $I$ varies. To make comparisons, we plot the effective $A_{\mu}$ charge $Q_{I}$ of topological excitations under PBC in the $z$-direction in Fig. 7(b1-b4). As is observed, $Q_{I}$ displays a translational symmetry, such that $Q_{I}=Q_{I+2}$. The comparison demonstrates that the random oscillation of $Q_{I}$ is due to the influence of boundary zero modes of $K$.
In contrast, for the (gapped) iCS field theories without Toeplitz braiding, even with OBC in $z$, $Q_{I}$ exhibits translational symmetry for $I$ corresponding to the $z$-bulk coordinates ($1\ll I\ll 2N$), such that $Q_{I}\simeq Q_{I+2}$, consistent with the charge fractionalization pattern observed under PBC. In Fig. 7(a5) and Fig. 7(b5), we present the fractionalized charges $Q_{I}$ of topological excitations in $K_{5}$-matrix iCS field theories. Fig. 7(a5) is under OBC while Fig. 7(b5) is under PBC. From this, we understand that in the absence of boundary zero modes of the $K$-matrix, (gapped) iCS field theories with PBC provide a good description to the $z$-bulk physics in OBC.